6 quarters,10 dimes ,6 nickles,and 6 pennes
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

6 quarters,10 dimes ,6 nickles,and 6 pennes
--------------
Do you have a question?

Answer by jim_thompson5910(35256)

Number of pennies = 6
Number of nickels = 6
Number of dimes = 10
Number of quarters = 6

Total value of pennies only: 6*0.01 = $0.06
Total value of nickels only: 6*0.05 = $0.30
Total value of dimes only: 10*0.1 = $1.00
Total value of quarters only: 6*0.25 = $1.50


Now add up the individual values:

0.06 + 0.30 + 1.00 + 1.50 = 2.86


So the total value of 6 pennies, 6 nickels, 10 dimes, and 6 quarters is $2.86